OpenStudy (anonymous):

((2p-2)/p)*7p^2/(4p-4)

OpenStudy (anonymous):

\[\frac{ 2p-2 }{p }*\frac{ 7p^2 }{ 4p-4 }\]

hartnn (hartnn):

2p-2 can be factored as 2 (p-1) factored out the 2 similarly, 4p-4 = 4 (p-1) \(\large \frac{ 2p-2 }{p }*\frac{ 7p^2 }{ 4p-4 }=\frac{ 2\cancel{(p-1)} }{\cancel{p} }*\frac{ 7p^\cancel{2} }{ 4\cancel{(p-1)} }\) and also,2/4 =1/2 you are left with 7p/2 only.